OpenStudy (rock_mit182):

Suppose that the functions and are defined as follows. u(x) = x^2 + 5 w(x) = \[\sqrt{x+3}\] find the following (w ° u)(1) (u ° w)(1)

OpenStudy (rock_mit182):

w(x^2 +5)= \[(\sqrt{x+3})^{2} +5\]

OpenStudy (rock_mit182):

(w ° u)(1) = x +8 = 9

OpenStudy (anonymous):

correct

OpenStudy (rock_mit182):

\[u (\sqrt{x+3}) =\] \[\sqrt{x ^{2}+5+3} =2x \sqrt{2}\]

OpenStudy (rock_mit182):

(u ° w )(1) = 2(1) sqroot(2)

OpenStudy (rock_mit182):

right ?